OpenStudy (anonymous):

suppose f(x)=x^2 what is the graph of f(x+2)

OpenStudy (campbell_st):

so if \[f(x) = x^2\] then \[f(x +2) = (x + 2)^2\] so the original function has be moved horizontally. the original graph has a zero at x = 0 by finding the zero of f(x +2) you'll have 1. the zero of the equation 2. the shift of the graph